Circular No. 5868

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA LUPI 1993 R. H. McNaught and G. J. Garradd, Anglo-Australian Observatory,

report a precise position of this apparent nova (cf. IAUC 5867) from an exposure obtained with the Uppsala Southern Schmidt telescope on Sept. 21.44 UT: α = 14^{h}31^{m}50^{s}.20, δ = −51^{o}10′33′′.1 (equinox 2000.0; uncertainty 0′′.3 in each coordinate). The corresponding 1950.0 position is α = 14^{h}28^{m}26^{s}.04, δ = −50^{o}57′18′′.2. Visual comparison of this exposure with the ESO B and R and SERC J surveys yields no candidate for this object to mag 17, and no stars in the immediate vicinity display obvious color.

CCD photometry by W. Liller, Viña del Mar, Chile: Sept. 19.987 UT, V = 7.97 ± 0.10, B–V = +0.18 ± 0.15; 20.994, V = 8.47 ± 0.02 (standard stars from nearby E5 region). A low-resolution objective prism spectrogram obtained with a 0.20-m reflector shows similarity to the spectra of other novae, including N Sgr 1993; the intensity of the Hα feature and nebular lines is at least twice the intensity of the surrounding continuum.

NOVA SAGITTARII 1993 Liller also reports the following additional CCD magnitudes: Sept.

17.058 UT, V = 8.51 ± 0.05; 18.015, 8.47; 19.041, 8.27; 19.987, 8.75. He further notes that, since Sept. 17.1, the intensity of the Hα feature plus neighboring nebular lines has increased from 3.0 to 4.2 times the continuum intensity.

Visual magnitude estimates, some forwarded by J. A. Mattei, AAVSO: Sept. 17.70 UT, 8.5 (L. Szentaskó, Budapest, Hungary); 18.14, 8.5 (P. Collins, Scottsdale, AZ); 18.792, 7.7 (J. Hers, Sedgefield, S. Africa); 18.8, 8.6 (D. Overbeek, Edenvale, S. Africa); 18.85, 8.2 (A. Pereira, Cabo da Roca, Portugal); 19.19, 8.2 (R. Royer, Lakewood, CA); 19.33, 8.1 (A. F. Jones, Nelson, New Zealand); 19.80, 8.4 (P. Schmeer, Bischmisheim, Germany); 19.85, 8.8 (Pereira); 20.02, 8.8 (M. Adams, Melbourne, FL); 20.05, 8.1 (C. Scovil, Stamford, CT); 20.06, 8.3 (J. Griese, Rocky Hill, CT).

COMET MUELLER (1993p) Total visual magnitude estimates: Aug. 24.64 UT, 13.8 (A. Nakamura,

Kuma, Japan, 0.60-m reflector); 27.65, 13.9 (Nakamura); Sept. 17.98, 13.4 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 19.18, 13.5 (C. S. Morris, Lockwood Valley, CA, 0.51-m reflector).

1993 September 21 Daniel W. E. Green